OpenStudy (anonymous):

If you flip a coin 3 times, what is the probability getting three tails??? (enter answer as a fraction in simplest terms.)

OpenStudy (anonymous):

1/8

OpenStudy (anonymous):

Use the distributive property to simplify the expression below: -3X+9/-3 Basiclly It Means -3x+9 Over -3 In Answer A.-x - 3 B.x + 9 C.x + 3 D.x - 3

OpenStudy (anonymous):

i know i sound stupid my my mind is drawing a blank

OpenStudy (anonymous):

I would list out the sample space: HHH HHT HTH THH HTT THT TTH TTT The only one with TTT is one of them out of 8 so: 1/8

OpenStudy (anonymous):

or \[\frac{1}{2}\times \frac{1}{2}\times \frac{1}{2}=\frac{1}{8}\]